OpenStudy (anonymous):

How do you get x + 2y + 4 = 0 into y=mx+b for a graph ??? then finding the 3 points

OpenStudy (anonymous):

Ok so solve for y

OpenStudy (anonymous):

so subtract x from both sides/?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so 2y + 4 = 1?

OpenStudy (anonymous):

what happened to your x...?

OpenStudy (anonymous):

2y - x + 4 = -1

OpenStudy (anonymous):

I don't understand what you're doing there

OpenStudy (anonymous):

where did you get -1 from?

OpenStudy (anonymous):

get y on its own

OpenStudy (anonymous):

0-1 = -1 isnt it?

OpenStudy (anonymous):

Isolate y

OpenStudy (anonymous):

\[x+2y+4 = 0\] \[2y+4 = -x\] \[2y = -x -4\] \[y = \frac{ -x-4 }{ 2 }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

No you're doing it wrong

OpenStudy (texaschic101):

x + 2y + 4 = 0 (subtract 4 from both sides) x + 2y + 4 - 4 = 0 - 4 (combine lime terms) x + 2y = -4 (now we want to subtract x from both sides) x + 2y - x = -x - 4 (combine like terms) 2y = -x - 4 (now divide both sides by 2) 2y/2 = -x/2 - 4/2 y = -1/2 - 2

OpenStudy (texaschic101):

batman got it right...I just went a step further